the wages of 12 men and 5 boys came to $446.40 per day. At a similar rate 10 men and 3 boys would earn $348.48 per day. Assuming each man and boy work an 8 hour day, find the hourly rate each is paid
Answer by ewatrrr(24785)


Hi
 Let x and y represent the hourly rate respectively
  12x + 5y = 446.40
  10x + 3y = 348.48 
 
  -7.2x -3y = -267.84    |multiply 1st EQ by -3/5 to eliminate y
  10x + 3y = 348.48 
          x =  80.64/2.8 = $28.80/hr is the hourly wage for the men
     and 10(28.80) + 3y = 348.48,  y = $20.16  the hourly wage for the boys

EQ 1:  $345.60 + $100.80= $446.40 checks  
Wish You the Best in your Studies.